find the shortest distance between the parallel lines with equations 5x-12y+33=0 and 5x-12y-6=0
Answer by edjones(8007)

5x-12y+33=0
5x-12y-6=0
.
5x-12y+33=0
-12y=-5x-33
y=5x/12 + 33/12
.
5x-12y-6=0
-12y=-5x+6
y=5x/12 - 1/2
.
The distance between the parallel lines is |33/12 + 1/2|=3.25
This is easier to see if we change the slope from 5x/12 (1st graph) to 0x/12 (2nd graph)

.
Ed
.

.